In 1988, bettys birthday was on a cold Wednesday in january. In 1993, on what day did her birthday occur?
A) Sunday B) Monday C) Tuesday D) Wednesday

Answer by Edwin McCravy(20064)


If a given date is on a certain day of the week one year, the following
year that date will advance one day of the week if there is (or was) no
February 29th between that date and and the next, and it will advance
two days if there is (or was) a February 29th between that date the first
year and that date the second year.

During the 5 years between that January Wednesday 1n 1988, and that date in
1993, there were two February 29ths, one in 1988 and another in 1992, so the
date advanced 1 day of the week 3 times and 2 days of the week 2 times, and
since 3 + 2*2 = 3+4 = 7, her birthday advanced 7 days from Wednesday, which
means her birthday was on Wednesday in 1993.  Choice D)

Edwin